# CalibRefine: Deep Learning-Based Online Automatic Targetless LiDAR–Camera Calibration with Iterative and Attention-Driven Post-Refinement

### v4

@@ Abstract
Accurate multi-sensor calibration is essential for deploying robust perception systems in applications such as autonomous driving and intelligent transportation. Existing LiDAR-camera calibration methods often rely on manually placed targets, preliminary parameter estimates, or intensive data preprocessing, limiting their scalability and adaptability in real-world settings. In this work, we propose a fully automatic, targetless, and online calibration framework, CalibRefine, which directly processes raw LiDAR point clouds and camera images. Our approach is divided into four stages: (1) a Common Feature Discriminator that trains on automatically detected objects–using relative positions, appearance embeddings, and semantic classes–to generate reliable LiDAR-camera correspondences, (2) a coarse homography-based calibration, (3) an iterative refinement to incrementally improve alignment as additional data frames become available, and (4) an attention-based refinement that addresses non-planar distortions by leveraging a Vision Transformer and cross-attention mechanisms. Through extensive experiments on two urban traffic datasets, we show that CalibRefine delivers high-precision calibration results with minimal human involvement, outperforming state-of-the-art targetless methods and remaining competitive with, or surpassing, manually tuned baselines. Our findings highlight how robust object-level feature matching, together with iterative and self-supervised attention-based adjustments, enables consistent sensor fusion in complex, real-world conditions without requiring ground-truth calibration matrices or elaborate data preprocessing. Code is available at https://github.com/radar-lab/Lidar_Camera_Automatic_Calibration
@@
LiDAR-camera calibration, extrinsic calibration, online automatic calibration, sensor fusion, traffic monitoring
@@ I Introduction
Reliable and accurate environment perception is crucial for applications such as autonomous driving, robotics, and intelligent transportation systems, enabling informed decisions and ensuring safe, efficient operations. However, single-sensor perception often encounters inherent limitations [1, 2]. Cameras provide rich visual detail but are sensitive to lighting variations and struggle with depth estimation, especially under poor illumination or dynamic conditions [3, 4]. LiDAR sensors, conversely, provide precise 3D geometric measurements robust to lighting changes, yet can be costly and suffer performance degradation under adverse weather [5, 6]. Consequently, multi-sensor fusion, particularly LiDAR-camera fusion, has gained prominence, integrating visual textures with accurate spatial data [7, 8]. Nonetheless, successful fusion fundamentally relies on accurate sensor calibration, as imprecise calibration severely compromises downstream perception accuracy [9, 10].
@@
Given these trade-offs, a fully automatic, targetless, and online calibration paradigm combines the most desirable attributes—removing cumbersome calibration objects, eliminating the need for human intervention, and adapting in real time to changing environments. However, despite its appeal, such a paradigm remains highly challenging, and existing approaches in this category exhibit various limitations. Motion-based methods often rely on additional hardware or specific sensor motion constraints, limiting real-world applicability [26, 27]. Hand-eye calibration, while classical, demands multiple accurate sensor poses, rendering it unsuitable for static or dynamically constrained setups [28]. Edge-based methods suffers from unreliable edge matching across modalities, since object boundaries differ significantly across LiDAR and camera data [29, 30, 5]. Mutual information-based methods are similarly unreliable due to varying LiDAR reflectance and camera illumination sensitivities [31, 32]. Recent deep learning–based methods directly regress calibration parameters (e.g., RegNet [23] and its variants [24, 1, 17, 33, 25, 34, 35]) but require initial manual calibrationsor projected depth maps, and suffer from limited generalization and computational overhead, making them less suitable for real-time applications [25, 17].
@@
To address these challenges, we propose CalibRefine, a fully automatic, targetless, online LiDAR–camera calibration framework. Our method directly processes raw LiDAR point clouds and camera images without initial calibration matrices or complex preprocessing. First, we leverage robust object detection algorithms—YOLOv8 for camera data and an octree-based DBSCAN approach for LiDAR—to identify individual objects. A novel Common Feature Discriminator then matches these cross-sensor object instances by learning relative positions, appearance embeddings (using ResNet [36] for camera images and PointNet++ [37] for LiDAR data), and semantic class information, forming reliable cross-sensor correspondences. Recognizing potential inaccuracies from initial matching, we further enhance calibration through two online refinement stages. The iterative refinement incrementally optimizes calibration using accumulated correspondence data, while the attention-based refinement employs a Vision Transformer [38] and cross-attention mechanisms to correct for non-planar distortions and depth variations, further improving calibration accuracy. Crucially, our framework bypasses the pitfalls of direct matrix regression and the need for projected LiDAR maps, and eliminates the reliance on heuristic preprocessing or manually labeled calibration matrices, offering a more principled, data-driven pipeline that is both computationally efficient and adaptable in real-time. By integrating domain-specific mature object detection methods, a reliable discriminator to identify cross-sensor correspondences, and dual-stage refinement, our approach bridges the existing research gap, achieving a stable and accurate LiDAR–camera calibration that is truly automatic, targetless, and online. Our contributions are summarized as follows:
@@
    Fully Automatic, Targetless, and Online Calibration Framework: We propose a novel calibration framework that directly processes raw LiDAR point clouds and camera images, eliminating the need for heuristic preprocessing, manually labeled calibration matrices, or initial calibration. This ensures generalizability and adaptability across diverse scenarios.
@@
Figure 2: Overall Structure of the Common Feature Discriminator.
@@ II Related Works
Extrinsic calibration methods for LiDAR–camera systems can generally be divided into two categories: target-based and targetless approaches. Target-based calibration relies on specially designed calibration targets and is thus commonly associated with manual, offline procedures. Moreover, these methods typically cannot handle real-time decalibrations, which are common in practical scenarios. In contrast, targetless methods extract features directly from natural scenes, making them well-suited for automatic, online calibration. These methods can be broadly categorized into motion-based, edge alignment-based, mutual information-based, and deep learning–based approaches.
@@ II-1 Motion-based Calibration
Motion-based calibration methods leverage sensor movements or relative poses derived from visual and LiDAR odometry to compute extrinsic parameters [39, 26, 10]. For instance, Petek et al. [39] utilize odometry paths from each sensor, aligning them through non-linear optimization and dense 2D–3D matching. Park et al. [26] similarly derive closed-form calibration from relative sensor transformations. Despite effectiveness under certain conditions, these methods heavily rely on accurate odometry or SLAM estimations, which are susceptible to noise and degenerate in scenarios with limited sensor motion (e.g., minimal rotation). Hand-eye calibration–based methods [28] further exacerbate this issue, requiring multiple precise sensor poses, thus limiting their practical applicability, particularly for static installations.
@@ II-2 Edge Alignment-based Calibration
Edge-based approaches attempt to align edges detected from LiDAR point clouds and camera images [5, 29, 30]. For example, Zhang et al. [5] transform the calibration problem into a cylindrical projection-based 2D–2D alignment task, while Li et al. [29] employ advanced edge extraction techniques such as the Segment Anything Model (SAM) combined with multi-frame filtering. However, reliably matching edges between sensors is inherently challenging due to modality differences—LiDAR captures sparse geometric structures, whereas camera edges reflect dense texture and lighting variations. These differences often lead to inaccurate feature correspondence and suboptimal calibration outcomes.
@@ II-3 Mutual Information-based Calibration
Figure 3: Schematic Overview of the Attention-based Refinement Process.
@@ II-4 Deep Learning-based Calibration
In summary, existing automatic, targetless, and online calibration methods commonly exhibit limitations including reliance on accurate odometry or sensor movement, difficulty in cross-modal edge matching, sensitivity to reflectance and illumination conditions, and computational inefficiencies. Furthermore, most existing methods fail to fully exploit the advances in object detection and feature extraction developed for LiDAR and camera data processing. To overcome these limitations, we propose CalibRefine, a fully automatic, targetless, and online calibration framework directly processing raw LiDAR point clouds and camera images, eliminating initial calibrations or elaborate preprocessing. Our approach integrates proven object detection algorithms and introduces a novel Common Feature Discriminator for robust cross-sensor correspondence matching. Furthermore, we employ a coarse-to-fine strategy combining iterative optimization and attention-driven refinement, enabling accurate and robust real-time calibration. By directly matching corresponding points across modalities, our approach facilitates a straightforward, one-shot, and end-to-end calibration process between the LiDAR and camera, significantly enhancing adaptability to real-world scenarios.
@@ III-A Problem Formulation
$\begin{bmatrix}u_{l}\\ v_{l}\\ 1\end{bmatrix}=\mathbf{H}\begin{bmatrix}x\\ y\\ 1\end{bmatrix}=\begin{bmatrix}h_{11}&h_{12}&h_{13}\\ h_{21}&h_{22}&h_{23}\\ h_{31}&h_{32}&h_{33}\end{bmatrix}\begin{bmatrix}x\\ y\\ 1\end{bmatrix},$ (1)
where $\hat{P_{l}}=(u_{l},v_{l})$ is the projection of a point $P_{l}=(x,y)$ in the LCS onto the camera image plane PCS. Notably, the objects or points on the 2D LiDAR plane and those on the camera image plane are derived from objects or points lying on a common plane (e.g., the ground plane), as shown in Fig. 5. This alignment justifies the use of 2D Homography for LiDAR–camera calibration, as it can be considered a planar homography induced by the common plane [43]. To solve for the Homography matrix, a set of $N$ points in the LCS and their corresponding points in the PCS is required. Although 4 points are theoretically sufficient, using more points allows optimization of the solution via a cost function that minimizes the geometric reprojection error [44, 42]. This error, which measures the alignment between $N$ projected LiDAR points ($\hat{P_{l}}=(\hat{u},\hat{v})$) and their corresponding 2D image pixel points ($P_{p}=(u,v)$), can be quantified either as the Average Euclidean Distance (AED)
$\begin{split}\mathcal{E}_{\text{AED}}&=\frac{1}{N}\sum_{i=1}^{N}\left\|P_{p}^{i}-\hat{P}_{l}^{i}\right\|_{2}\\ &=\frac{1}{N}\sum_{i=1}^{N}\sqrt{(u^{i}-\hat{u}^{i})^{2}+(v^{i}-\hat{v}^{i})^{2}},\end{split}$ (2)
or as the Root Mean Square Error (RMSE)
$\begin{split}\mathcal{E}_{\text{RMSE}}&=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left\|P_{p}^{i}-\hat{P}_{l}^{i}\right\|_{2}^{2}}\\ &=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left[(u^{i}-\hat{u}^{i})^{2}+(v^{i}-\hat{u}^{i})^{2}\right]}.\end{split}$ (3)
@@ III-C Common Feature Discriminator
    Reflection Intensity: LiDAR measures reflection intensity based on surface material properties, while cameras capture similar information through brightness and contrast.
@@
To achieve this, we propose the Common Feature Discriminator, a deep learning–based model that leverages advanced feature extraction to learn and extract shared features from LiDAR and camera data, thereby enabling effective object matching and correspondence identification. The first step is to detect and crop individual objects from each sensor’s output. For camera-based object detection, we adopt YOLOv8 [46] to robustly detect objects in images and generate bounding boxes around them. In parallel, LiDAR-based object detection is performed using an octree-based change detection algorithm [47] followed by DBSCAN clustering, which segments the point cloud into clusters, each hypothesized to belong to a distinct object. Since the LiDAR and camera frames are time-synchronized, each LiDAR cluster and corresponding camera bounding box at the same timestamp can be treated as candidate detections from complementary modalities.
@@
By jointly analyzing $\Delta\mathbf{p}$, $\mathbf{z}_{L}$, $\mathbf{z}_{C}$, and $\hat{c}_{L},\hat{c}_{C}$, the Common Feature Discriminator robustly determines whether the LiDAR and camera detections refer to the same underlying object, even when the modalities present substantially different raw representations, thereby enabling the system to automatically match and associate objects across LiDAR and camera views. This module, integrated with LiDAR and camera object detectors, constitutes the foundation of an end-to-end cross-sensor object matching workflow. Specifically, time-synchronized LiDAR and camera frames are processed in parallel, and bounding boxes (camera) or point clusters (LiDAR) are cropped and fed into the discriminator to obtain pairwise correspondence labels. The resulting high-confidence matches form the cornerstone for computing the extrinsic calibration matrix that aligns the LiDAR and camera coordinate frames.
@@ III-D Homography-based Calibration Matrix Estimation
Algorithm 1 Iterative LiDAR–Camera Calibration Refinement
@@ III-E Iterative Refinement Process
By systematically incorporating additional correspondences at each iteration, this optimization loop converges toward a more robust calibration matrix. It maintains the practical advantages of the initial deep learning–based matching while progressively enhancing accuracy through redundancy and extended spatial coverage. Moreover, its iterative nature naturally accommodates runtime changes in the environment, thus helping to mitigate potential decalibration over long-term operation. Notably, we opt for a greedy bipartite matching [48] approach rather than the more popular Hungarian algorithm for several practical reasons. Ideally, each LiDAR detection would correspond to exactly one camera detection, and bipartite graph matching would produce a one-to-one mapping that minimizes the overall matching cost. However, real-world conditions deviate from this ideal scenario: variations in field of view and detection capabilities can lead to certain objects being detected by only one sensor. For example, LiDAR may capture distant objects outside the camera’s range, whereas a camera may pick up small or reflective objects that the LiDAR cannot reliably detect. Given these discrepancies, the goal of bipartite graph matching is to identify the best subset of matching pairs, without forcing all detections from both sensors to be paired. Greedy bipartite matching is well-suited to this task, as it prioritizes finding and accumulating the lowest-cost matches while allowing some objects to remain unmatched if no suitable pair exists. In contrast, the Hungarian algorithm aims for an optimal, one-to-one, and complete assignment—i.e., pairing every detection from both sensors—an assumption that does not hold in many real-world LiDAR–camera detections. Such forced one-to-one pairings can degrade matching quality when unmatchable objects are forced to pair with unrelated detections.

### v5

@@ Abstract
Accurate multi-sensor calibration is essential for deploying robust perception systems in applications such as autonomous driving and intelligent transportation. Existing LiDAR-camera calibration methods often rely on manually placed targets, preliminary parameter estimates, or intensive data preprocessing, limiting their scalability and adaptability in real-world settings. In this work, we propose a fully automatic, targetless, and online calibration framework, CalibRefine, which directly processes raw LiDAR point clouds and camera images. Our approach is divided into four stages: (1) a Common Feature Discriminator that leverages relative spatial positions, visual appearance embeddings, and semantic class cues to identify and generate reliable LiDAR-camera correspondences, (2) a coarse homography-based calibration that uses the matched feature correspondences to estimate an initial transformation between the LiDAR and camera frames, serving as the foundation for further refinement, (3) an iterative refinement to incrementally improve alignment as additional data frames become available, and (4) an attention-based refinement that addresses non-planar distortions by leveraging a Vision Transformer and cross-attention mechanisms. Extensive experiments on two urban traffic datasets demonstrate that CalibRefine achieves high-precision calibration with minimal human input, outperforming state-of-the-art targetless methods and matching or surpassing manually tuned baselines. Our results show that robust object-level feature matching, combined with iterative refinement and self-supervised attention-based refinement, enables reliable sensor alignment in complex real-world conditions without ground-truth matrices or elaborate preprocessing. Code is available at https://github.com/radar-lab/Lidar_Camera_Automatic_Calibration
@@
LiDAR-camera calibration, extrinsic calibration, online automatic calibration, sensor fusion, deep learning
@@ I Introduction
Reliable and accurate environment perception is crucial for applications such as autonomous driving, robotics, and intelligent transportation systems, enabling informed decisions and ensuring safe, efficient operations. However, single-sensor perception often encounters inherent limitations [1, 2, 3]. Cameras provide rich visual detail but are sensitive to lighting variations and struggle with depth estimation, especially under poor illumination conditions [4, 5]. LiDAR sensors, conversely, provide precise 3D geometric measurements robust to lighting changes [6], yet can be costly and suffer performance degradation under adverse weather. Consequently, multi-sensor fusion, particularly LiDAR-camera fusion, has gained prominence, integrating visual textures with accurate spatial data [7, 8]. Nonetheless, successful fusion fundamentally relies on accurate sensor calibration, as imprecise calibration severely compromises downstream perception accuracy [9, 10].
@@
Given these trade-offs, a fully automatic, targetless, and online calibration paradigm combines the most desirable attributes—removing cumbersome calibration objects, eliminating the need for human intervention, and adapting in real time to changing environments. However, despite its appeal, such a paradigm remains highly challenging, and existing approaches in this category exhibit various limitations. Motion-based methods often rely on additional hardware or specific sensor motion constraints, limiting real-world applicability [26]. Hand-eye calibration, while classical, demands multiple accurate sensor poses, rendering it unsuitable for static or dynamically constrained setups [27]. Edge-based methods suffer from unreliable edge matching across modalities, since object boundaries differ significantly across LiDAR and camera data [28, 29, 16]. Mutual information-based methods are similarly unreliable due to varying LiDAR reflectance and camera illumination sensitivities [30, 31]. Recent deep learning–based methods directly regress calibration parameters (e.g., RegNet [23] and its variants [24, 2, 17, 32, 33, 34]) but require initial manual calibrations or projected depth maps, and suffer from limited generalization and computational overhead, making them less suitable for real-time applications [17].
@@
To address these challenges, we propose CalibRefine, a fully automatic, targetless, online LiDAR–camera calibration framework. Our method directly processes raw LiDAR point clouds and camera images without initial calibration matrices or complex preprocessing. First, we leverage established object detection algorithms—YOLOv8 for camera data and an octree-based DBSCAN approach for LiDAR—to identify individual objects. A novel Common Feature Discriminator then matches these cross-sensor object instances by learning relative positions, appearance embeddings (using ResNet [35] for camera images and PointNet++ [36] for LiDAR data), and semantic class information, forming reliable cross-sensor correspondences. Recognizing potential inaccuracies from initial matching, we further enhance calibration through two online refinement stages. The iterative refinement incrementally optimizes calibration using accumulated correspondence data, while the attention-based refinement employs a Vision Transformer [37] and cross-attention mechanisms to correct for non-planar distortions and depth variations, further improving calibration accuracy. Crucially, our framework bypasses the pitfalls of direct matrix regression and the need for projected LiDAR maps, and eliminates the reliance on heuristic preprocessing or manually labeled calibration matrices, offering a more principled, data-driven pipeline that is both computationally efficient and adaptable in real-time. By integrating domain-specific mature object detection methods, a reliable discriminator to identify cross-sensor correspondences, and dual-stage refinement, our approach bridges the existing research gap, achieving a stable and accurate LiDAR–camera calibration that is truly automatic, targetless, and online. Our contributions are summarized as follows:
@@
    Fully Automatic, Targetless, and Online Calibration Framework: We propose a novel calibration framework that directly processes raw LiDAR point clouds and camera images, eliminating the need for heuristic preprocessing, manually labeled calibration matrices, or prior calibration. This ensures generalizability and adaptability across diverse scenarios.
@@
Figure 2: Key Components of Our Proposed Method: (a) Common Feature Discriminator, (b) Attention-based Refinement Module.
@@ II Related Works
Extrinsic calibration methods for LiDAR–camera systems can generally be divided into two categories: target-based and targetless approaches. Target-based calibration [38, 39] relies on specially designed calibration targets and is thus commonly associated with manual, offline procedures. Moreover, these methods typically cannot handle real-time decalibrations, which are common in practical scenarios. In contrast, targetless methods extract features directly from natural scenes, making them well-suited for automatic, online calibration. These methods can be broadly categorized into motion-based, edge alignment-based, mutual information-based, and deep learning–based approaches.
@@ II-1 Motion-based Calibration
Motion-based calibration methods leverage sensor movements or relative poses derived from visual and LiDAR odometry to compute extrinsic parameters [40, 41, 10, 27]. For instance, Petek et al. [40] utilize odometry paths from each sensor, aligning them through non-linear optimization and dense 2D–3D matching. Zhang et al. [41] similarly derive the calibration from relative sensor SLAM trajectories. Despite effectiveness under certain conditions, these methods heavily rely on accurate odometry or SLAM estimations, which are susceptible to noise and degenerate in scenarios with limited sensor motion (e.g., minimal rotation). Hand-eye calibration–based methods [27] further exacerbate this issue, requiring multiple precise sensor poses, thus limiting their practical applicability, particularly for static installations.
@@ II-2 Edge Alignment-based Calibration
Edge-based approaches attempt to align edges detected from LiDAR point clouds and camera images [16, 28, 29]. For example, Zhu et al. [29] extract edge features of LiDAR-Camera data for edge alignment based solely on laser reflectivity, while Li et al. [28] employ advanced edge extraction techniques such as the Segment Anything Model (SAM) combined with multi-frame filtering. However, reliably matching edges between sensors is inherently challenging due to modality differences—LiDAR captures sparse geometric structures, whereas camera edges reflect dense texture and lighting variations. These differences often lead to inaccurate feature correspondence and suboptimal calibration outcomes.
@@ II-4 Deep Learning-based Calibration
In summary, existing automatic, targetless, and online calibration methods commonly exhibit limitations including reliance on accurate odometry or sensor movement, difficulty in cross-modal edge matching, sensitivity to reflectance and illumination conditions, and computational inefficiencies. Furthermore, most existing methods fail to fully exploit the advances in object detection and feature extraction developed individually for LiDAR and camera data processing. To overcome these limitations, we propose CalibRefine, a fully automatic, targetless, and online calibration framework that directly processes raw LiDAR point clouds and camera images, eliminating initial calibrations or elaborate preprocessing. Our approach integrates proven object detection algorithms and introduces a novel Common Feature Discriminator for robust cross-sensor correspondence matching. Furthermore, we employ a coarse-to-fine strategy combining iterative optimization and attention-driven refinement, enabling accurate and robust real-time calibration. By directly matching corresponding points across modalities, our approach facilitates a straightforward, one-shot, and end-to-end calibration process between the LiDAR and camera, significantly enhancing adaptability to real-world scenarios.
@@ III-A Problem Formulation
$\begin{bmatrix}\hat{u}\\ \hat{v}\\ 1\end{bmatrix}=\mathbf{H}\begin{bmatrix}x\\ y\\ 1\end{bmatrix}=\begin{bmatrix}h_{11}&h_{12}&h_{13}\\ h_{21}&h_{22}&h_{23}\\ h_{31}&h_{32}&h_{33}\end{bmatrix}\begin{bmatrix}x\\ y\\ 1\end{bmatrix},$ (1)
where ($\hat{P_{l}}=(\hat{u},\hat{v})$) is the projection of a point $P_{l}=(x,y)$ in the LCS onto the camera image plane PCS. Notably, the objects or points on the 2D LiDAR plane and those on the camera image plane are derived from objects or points lying on a common plane (e.g., the ground plane), as shown in Fig. 4. This alignment justifies the use of 2D Homography for LiDAR–camera calibration, as it can be considered a planar homography induced by the common plane [45]. To solve for the Homography matrix, a set of $N$ points in the LCS and their corresponding points in the PCS is required. Although 4 points are theoretically sufficient, using more points allows optimization of the solution via a cost function that minimizes the geometric reprojection error [46, 44]. This error, which measures the alignment between $N$ projected LiDAR points and their corresponding 2D image pixel points ($P_{p}=(u,v)$), can be quantified either as the Average Euclidean Distance (AED) or as the Root Mean Square Error (RMSE):
$\mathcal{E}_{\text{AED}}=\frac{1}{N}\sum_{i=1}^{N}\left\|P_{p}^{i}-\hat{P}_{l}^{i}\right\|_{2}=\frac{1}{N}\sum_{i=1}^{N}\sqrt{(u^{i}-\hat{u}^{i})^{2}+(v^{i}-\hat{v}^{i})^{2}}$ (2)
$\mathcal{E}_{\text{RMSE}}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left\|P_{p}^{i}-\hat{P}_{l}^{i}\right\|_{2}^{2}}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left[(u^{i}-\hat{u}^{i})^{2}+(v^{i}-\hat{v}^{i})^{2}\right]}$ (3)
@@ III-C Common Feature Discriminator
    Reflection Intensity: LiDAR measures reflection intensity based on surface properties, while cameras capture similar information through brightness and contrast.
@@
To achieve this, we propose the Common Feature Discriminator (see Fig. 2(a)), a deep learning–based model that leverages advanced feature extraction to learn and extract shared features from LiDAR and camera data, thereby enabling effective object matching and correspondence identification. The first step is to detect and crop individual objects from each sensor’s output. For camera-based object detection, we adopt YOLOv8 [48] to robustly detect objects in images and generate bounding boxes around them. In parallel, LiDAR-based object detection is performed using an octree-based change detection algorithm [49] followed by DBSCAN clustering, which segments the point cloud into clusters, each hypothesized to belong to a distinct object. Since the LiDAR and camera frames are time-synchronized, each LiDAR cluster and corresponding camera bounding box at the same timestamp can be treated as candidate detections from complementary modalities.
@@
By jointly analyzing $\Delta\mathbf{p}$, $\mathbf{z}_{L}$, $\mathbf{z}_{C}$, and $\hat{c}_{L},\hat{c}_{C}$, the Common Feature Discriminator robustly determines whether the LiDAR and camera detections refer to the same underlying object, even when the modalities present substantially different raw representations, thereby enabling the system to automatically match and associate objects across LiDAR and camera views. This module, integrated with LiDAR and camera object detectors, constitutes the foundation of an end-to-end cross-sensor object matching workflow. Specifically, time-synchronized LiDAR and camera frames are processed in parallel, and bounding boxes or point clusters are cropped and fed into the discriminator to obtain pairwise correspondence labels. The resulting high-confidence matches form the cornerstone for computing the extrinsic calibration matrix that aligns the LiDAR and camera coordinate frames.
@@ III-D Homography-based Calibration Matrix Estimation
Algorithm 1 Iterative Refined Calibration
@@ III-E Iterative Refinement Process
By systematically incorporating additional correspondences at each iteration, this optimization loop converges toward a more robust calibration matrix. It maintains the practical advantages of the initial deep learning–based matching while progressively enhancing accuracy through redundancy and extended spatial coverage. Moreover, its iterative nature naturally accommodates runtime changes in the environment, thus helping to mitigate potential decalibration over long-term operation. Notably, we opt for a greedy bipartite matching [50] approach rather than the more popular Hungarian algorithm for several practical reasons. Ideally, each LiDAR detection would correspond to exactly one camera detection, and bipartite graph matching would produce a one-to-one mapping that minimizes the overall matching cost. However, real-world conditions deviate from this ideal scenario: variations in the field of view and detection capabilities can lead to certain objects being detected by only one sensor. For example, LiDAR may capture distant objects outside the camera’s range, whereas a camera may pick up small or reflective objects that the LiDAR cannot reliably detect. Given these discrepancies, the goal of bipartite graph matching is to identify the best subset of matching pairs, without forcing all detections from both sensors to be paired. Greedy bipartite matching is well-suited to this task, as it prioritizes finding and accumulating the lowest-cost matches while allowing some objects to remain unmatched if no suitable pair exists. In contrast, the Hungarian algorithm aims for an optimal, one-to-one, and complete assignment—i.e., pairing every detection from both sensors—an assumption that does not hold in many real-world LiDAR–camera detections. Such forced one-to-one pairings can degrade matching quality when unmatchable objects are forced to pair with unrelated detections.